Database last updated on Sun, 01 Feb 2026 09:10:52 +0000 by localhost
Critical warnings from b-files' comments for days between 2023340 and 2023340 (0 sec, 4 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| hp: | Spectrometer alignment trouble. | 21 | 1 | | | #085 | 14:08:47 | | 1 | -1 |
| hg: | Lamp not detected f1= 0 f4= 47 | 2 | 1 | | | #085 | 16:01:22 | | 1 | -1 |
| hg: | Lamp not on... hg test terminated F1= 0 F4= 47 | 2 | 1 | | | #085 | 16:01:22 | | 1 | -1 |
| re: | Azimuth tracker ZEROING FAILURE | 1 | 2 | | | #085 | 18:36:15 | | 1 | -1 |
Instrument constant changes for the week before 2023340 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 16 rows)
Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)
| Serial | YYYYJJJ | HG large | max correction |
| 085 | 2023340 | 112 | 1051.0 |
HP with step corrections larger than 20 steps (0 sec, 1 rows)
| Serial | YYYYJJJ | HP large | max correction |
| 085 | 2023340 | 5 | 79.9 |
Non-critical warnings from b-files' comments for days between 2023340 and 2023340 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| re: | DEL key pressed | 4 | 99 | | | #085 | 17:10:16 | | 1 | -1 |
| pd: | Going to menu | 4 | 99 | | | #085 | 17:21:55 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | hg | hp |
| 340 | TORONTO | #085 | 31 | 5 |
(0 sec)
Dead time (0 sec)
Nothing to report.
Standard lamp R6 (1 sec)
Nothing to report.
Standard lamp R5 (1 sec)
Nothing to report.
Number of days since last dome/window cleaning (0 sec, 3 rows)
| serial | days |
| 085 | more than a year |
| 8 | more than a year |
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 085 | 2023338 | 002 |
Suspicious HG test results (correlation<0.98 or counts<5000). (0 sec, 1 rows)
| Serial | Jday | count | Peak | min.corr. | max.corr. | 30-day history |
| #085 | 340 | 28 | 150048 | 0 | 0.9796 |